Solve the differential Squation. 1 + 4 dt ut? + 14t2 du = +++ +C X 3 t
Answer #1

Please see the solution and if you have any doubt please comme, provide whole equation (solution) in the box

Thanks

Solution dy dt 1+14 ut tut t du dt (1+19) ť (utu) (ut u du = (1+ta) dt Scut u du = |( + x²) dt 3 U² us j + c L + 5 2 3 u² I t